OpenStudy (anonymous):

Maria looked at the architectural plan of a room with four walls in which the walls meet each other at right angles. The length of one wall in the plan is 13 inches. The length of the diagonal of the floor of the room in the plan is approximately 15.26 inches. Is the room in the shape of a square? Explain how you determined your answer. Show all your work. I'm pretty sure it's not a square, would i do 13 on both sides, then 13 squared twice add them and find the square root? that's how I got my answer but can you confirm thats the way to do it?

OpenStudy (anonymous):

i got 18.38 and so on

OpenStudy (sepeario):

I think you are on the right side. I think what it's trying to test is your application of Pythagoras' Theorem.

OpenStudy (sepeario):



OpenStudy (sepeario):

So what we have to have to do here is apply Pythagoras to that right angled triangle. Do you know how to do that?

OpenStudy (anonymous):

First since one side was 13 I made the other side 13. Then i multiplied 13 by 13 and got 169, i did the same thing to the other side and 169+169=338. The square root of 338 is 18.3847763109, therefore it isnt a square because it's not equal to 15.26 Yes its Pythagorean

OpenStudy (sepeario):

Yea that's pretty smart actually. There is another way to do it.

OpenStudy (anonymous):

a squared+ b squared = c squared ?

OpenStudy (sepeario):

You could instead say \[15.26^2-13^2=x^2\]

OpenStudy (sepeario):

Yes so I think you understand it pretty well. Nice!

OpenStudy (anonymous):

thanks for your help!

OpenStudy (sepeario):

all good.